Status Report

ISS On-Orbit Status 4 Jan 2003

By SpaceRef Editor
January 4, 2003
Filed under , ,
ISS On-Orbit Status 4 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.

A rest day for the crew, except for the regular weekend tasks.

FE-1 Nikolai Budarin started his Saturday off by completing the daily checkup of the BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) experiment for the research of plant growth and development under spaceflight conditions.

Budarin then performed the periodic inspection of the BRPK condensate water separator.

CDR Ken Bowersox, Budarin and FE-2/SO Don Pettit joined in the weekly 3-hr. housecleaning of their abode, which includes removal of food waste products, periodic cleaning of compartments with vacuum cleaner, and wet cleaning of surfaces.

Pettit took the daily ACS (atmosphere control and supply) system CO2 (carbon dioxide) readout with the CDMK (CO2 monitoring kit) as part of the current measurement program to help resolve discrepancies between ppCO2 (carbon dioxide partial pressure) readings by the SM gas analyzer and U.S. MCA (major constituents analyzer).

Following up on yesterday’s teleconference, the crew today discussed the upcoming Stage EVA (extravehicular activity) with the ground via S-band, focusing on timeline and procedures.  The EVA is tentatively scheduled for 1/15, i.e., well before Progress 10P arrival with the urgently awaited MSG (microgravity science glovebox) replacement parts.  [Major tasks on the timeline for the 6-hr. EVA are: (1) configure SFU (squib firing unit) on P1 to launch condition; (2) release 10 launch restraint locks (5 zenith, 5 nadir) on the P1 radiator beam; (3) inspect radiator beam and monitor its deployment; (4) reconfigure SFU; (5) clean Node nadir CBM (common berthing module); (6) install a tool box on CETA (crew and equipment translation aid) cart #2; (7) relocate and reconfigure CETA foot restraints; (8) install a light fixture on CETA #2; and (9) perform get-ahead tasks as open time permits.  It will be the 50th EVA for ISS assembly, with 25 from Shuttle and 25 from station (16 of these from the Joint Airlock).]

Regarding yesterday’s ZCG (Zeolite crystal growth) payload work, the furnace is heating up as expected and “all looks great”, to the joy of the ground teams.  Bowersox was thanked for successfully changing out the hard drives to correct a timing problem in the software.    [The timing problem was fixed. Although the back-up CPU (central processor unit) was not recovered; the ZCG furnace operations continued on the primary CPU.  Ground commanding to mix the autoclaves went well with the exception of one autoclave, tube 4, which appeared to jam. This autoclave malfunctioned due to the particular sand-like nature of the solution. The manual activation performed by Sox went well and saved a sample that would have otherwise been lost.  Sox’s judgments were invaluable to the ZCG team and the success of the ZCG experiment is directly attributable to his actions. The activation of these spare autoclaves will increase science output by more than 25%.]

The daily routine servicing task of SOSh life support systems maintenance (incl. ASU toilet system) was completed by Budarin, while Bowersox performed the regular daily status checkup of Lab payloads (PCG-STES-10, ZCG), and Don Pettit prepared the daily IMS inventory update file for downlinking.

Don downlinked information regarding HRM (heart rate monitor) transmitter serial numbers.  As he had suspected, the transmitter he is using was left over from Expedition 4 (E4), and the difficulty he had with the transmission of heart rate signals is believed to be due to a weak battery in the E4 transmitter  He was advised to start using the E6 spare and to conduct, along with Sox, a test of this theory during their CEVIS sessions.

All crewmembers completed their daily physical exercise program on TVIS, CEVIS, and RED.

SpaceRef staff editor.